### v2

# On the Source Model Key Agreement Problem

Abstract
We consider the source model key agreement problem involving two legitimate parties and an eavesdropper who observe $n$ i.i.d. samples of $X$ and $Y$ and $Z$ respectively. In this paper, we focus on one of the simplest instances where the key capacity remains open, specifically when $X$ and $Y$ are binary random variables and $Z$ is a function of the pair $(X,Y)$. The best-known upper bound on the key capacity is characterized by an inf-max optimization problem that generally lacks a closed-form solution. We provide general conditions under which the upper bound reduces to $I(X;Y)$. As an example, we consider the XOR setting in which $X$ and $Y$ are binary, and $Z$ is the XOR of $X$ and $Y$. The upper bound reduces to $I(X;Y)$ for this source. Next, we conjecture that the rate $I(X;Y)$ is not achievable for the XOR source, and provide some ideas that might be useful for developing a new upper bound on the source model problem.
1 Introduction
The two-party source model problem consists of two legitimate parties, Alice and Bob, who aim to establish a shared secret key hidden from an eavesdropper, Eve [1, 2]. Alice, Bob, and Eve observe $n$ i.i.d. samples of $X$ and $Y$ and $Z$ respectively. Random variables $X,Y$ and $Z$ are distributed according to $P_{X,Y,Z}$, known to all parties. Through authenticated but public communication, Alice and Bob collaborate to generate a key that remains secret from Eve (extensions to multiple parties and continuous random variables can be found in [3, 4, 5, 6, 7, 8]; please see [9] for a detailed treatment). In a non-interactive setting, Alice and Bob locally generate messages $F_{1}$ and $F_{2}$ from $P_{F_{1}|X^{n}}$ and $P_{F_{2}|Y^{n}}$, respectively, and exchange these messages. Then Alice and Bob create their own key from $K_{A}\sim P_{K_{A}|X^{n},F_{1},F_{2}}$ and $K_{B}\sim P_{K_{B}|Y^{n},F_{1},F_{2}}$ respectively. On the other hand, in the interactive setting, Alice initiates the process by generating a public message $F_{1}$ using the distribution $P_{F_{1}|X^{n}}$ and shares it with Bob. Bob then responds with $F_{2}$ derived from $P_{F_{2}|Y^{n},F_{1}}$. This exchange continues for $r$ rounds, during which both senders alternately generate and share public messages $F_{3},F_{4},\cdots F_{r}$. At the end of these interactions, Alice derives a key $K_{A}$ based on $P_{K_{A}|X^{n},F_{1:r}}$ while Bob produces a corresponding key $K_{B}$ using $P_{K_{B}|Y^{n},F_{1:r}}$. In an $(n,R,\delta)$ code, the keys must match with probability $1-\delta$:
$\mathbb{P}[K_{A}=K_{B}]\geq 1-\delta,$ (1)
where $\delta$ is a small positive value. The key rate is defined as $R=\frac{1}{n}H(K_{A})$. $R$ must exhibit high entropy, quantified as:
$R\geq\frac{1}{n}\log(|\mathcal{K}|)-\delta,$ (2)
where $\mathcal{K}$ denotes the alphabet of $K_{A}$ and $K_{B}$. Furthermore, the keys should be nearly independent of Eve’s observations, formalized by:
$\frac{1}{n}I(K_{A};Z^{n},F_{1:r})\leq\delta.$ (3)
Note that the value of $r$ for non-interactive regime is $2$. We note that if we omit the factor $1/n$ in equation (3), we obtain a seemingly stronger notion of secrecy; however, this is equivalent to the weaker form presented in (3). For a proof of equivalence, see [10].
We say that the rate $R$ is achievable if for every $\delta>0$, there is an $(n,\delta,R)$ code. The maximum achievable key rate is known as the source model secret-key (SK) capacity. It is denoted by $S(X;Y\|Z)$ in the interactive setting, and by $S_{\text{ni}}(X;Y\|Z)$ in the non-interactive setting.
Finding the secret key capacity remains an open problem in general scenarios, including both the interactive and non-interactive communication settings. Conditions that imply $S(X;Y\|Z)>0$ are studied in [2, 11, 12, 13]. The exact formula for secret key capacity is known only for some instances, such as when $X,Y$ and $Z$ form a Markov chain in some order. As a side result, we derive the capacity for a new class of sources, which we call “deterministic-erasure” sources in Appendix A.
One scenario where the capacity is known occurs when communication is restricted to one-way transmission from one party to another. The one-way SK capacity from $X$ to $Y$ equals [1]:
$S_{\text{ow}}(X\rightarrow Y\|Z)=\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}I(U;Y|V)-I(U;Z|V),$ (4)
where the conditions $|\mathcal{V}|\leq|\mathcal{X}|$ and $|\mathcal{U}|\leq|\mathcal{X}|$ can be imposed on $\mathcal{V}$, $\mathcal{U}$, and $\mathcal{X}$ representing the alphabets associated with $V$, $U$, and $X$, respectively [9]. [^1] We can equivalently express $S_{\text{ow}}(X\rightarrow Y\|Z)=\max_{V-\!\!\!\!\minuso\!\!\!\!-U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}I(U;Y|V)-I(U;Z|V),$ (5) This follows from noting that changing $U$ to $(U,V)$ will not change the mutual information terms, while introducing the additional Markov chain constraint. However, this comes at the cost of a larger cardinality bound on $U$. It is evident that $S_{\text{ow}}(X\rightarrow Y\|Z)$ establishes a lower bound on the SK capacity $S(X;Y\|Z)$.
An early upper bound on the key capacity was [2, 1]
$S(X;Y\|Z)\leq I(X;Y|Z).$
Authors in [1, pp. 1126, Remark 2],[11] improved this upper bound by realizing that degrading Eve cannot decrease the key rate, i.e., for any degradation $J$ of $Z$ through a Markov chain $J-\!\!\!\!\minuso\!\!\!\!-Z-\!\!\!\!\minuso\!\!\!\!-(X,Y)$ we have
$S(X;Y\|Z)\leq S(X;Y\|J)\leq I(X;Y|J).$
The quantity
$\min_{J-\!\!\!\!\minuso\!\!\!\!-Z-\!\!\!\!\minuso\!\!\!\!-(X,Y)}I(X;Y|J)$
is called the intrinsic conditional mutual information. More generally, we can relax the degraded assumption $J-\!\!\!\!\minuso\!\!\!\!-Z-\!\!\!\!\minuso\!\!\!\!-(X,Y)$ to a less noisy condition. The inequality
$S(X;Y\|Z)\leq S(X;Y\|J)$
holds if the channel $P_{Z|X,Y}$ is less noisy than the channel $P_{J|X,Y}$. Even more generally, for any arbitrary $P_{J|X,Y}$, we have [14]
$S(X;Y\|Z)-S(X;Y\|J)\leq\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(U;J|V)-I(U;Z|V)=S_{\text{ow}}(XY\rightarrow Z|J)$ (6)
Observe that the right hand side of the above equation vanishes when $P_{Z|X,Y}$ is less noisy than the channel $P_{J|X,Y}$, and was interpreted as “the penalty of deviating from the less-noisy condition” in [13]. Using (6) and the upper bound
$S(X;Y|J)\leq I(X;Y|J)$
the following upper bound was proposed in [14]:
$S(X;Y|Z)\leq\Psi(X;Y\|Z)$ $\triangleq\inf_{J}\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(X;Y|J)+I(U;J|V)-I(U;Z|V)$ (7)
Note that $J$ serves as an auxiliary receiver in the sense of [15]. The above bound is the best known upper on the key capacity. It involves a min-max optimization problem, which makes computing the upper bound challenging. In fact, it is not known whether the infimum over $J$ in (7) is a minimum (see [14] for a discussion on the computability of this bound). A weaker, but more convenient upper bound was also proposed in [14]
Proposition 1 ([14]).
For any $P_{X,Y,Z}$ we have:
$S(X;Y\|Z)$ $\leq\inf_{J}\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(X;Y|J)+I(U;J|V)-I(U;Z|V)$ (8)
$\leq\min_{J}I(X;Y|J)+I(X,Y;J|Z).$ (9)
We denote
$\hat{\Psi}(X;Y\|Z)\triangleq\min_{J}I(X;Y|J)+I(X,Y;J|Z).$
Notice that the upper bound $\Psi(X;Y\|Z)$ in (7) depends solely on $P_{X,Y,J}$ and $P_{X,Y,Z}$, meaning the minimization is restricted to $P_{J|X,Y}$. However the weaker upper bound $\hat{\Psi}(X;Y\|Z)$ in (9) requires minimization over $P_{J|X,Y,Z}$. On the other hand, the weaker upper bound $\hat{\Psi}(X;Y\|Z)$ does not involve the auxiliary random variables $U$ and $V$.
The entire research presented in this paper is motivated by our desire to study $S(X;Y\|Z)$ in what we believe is the simplest case where the key capacity remains unknown, namely, when $X$ and $Y$ are binary random variables, and $Z$ is a function of the pair $(X,Y)$. The main results of this paper are as follows:
  • We observe that $\hat{\Psi}(X;Y\|Z)=\Psi(X;Y\|Z)$ if $Z$ is a function of $(X,Y)$. This is immediate by the choice of $U=X,Y$ and $V=Z$ in (8) which would reduce the expression to that in (9).
  • We find sufficient conditions provided in Theorem 1 and Lemma 1 that imply that
    $\hat{\Psi}(X;Y\|Z)=\Psi(X;Y\|Z)=I(X;Y).$
    In particular, the sufficient conditions are satisfied when $X$ and $Y$ are binary random variables (with an arbitrary joint distribution) and $Z$ is the XOR of $X$ and $Y$ (see Theorem 2). We call this setting the XOR case. We also discuss other binary operations such as AND or OR in Theorem 2.
  • We conjecture that $S(X;Y\|Z)\neq I(X;Y)$ in the XOR case (see Section 4). Showing this would require developing a new upper bound for the source model problem. Partial results are also provided in our attempt to develop a new upper bound.
The rest of this paper is organized as follows: Section 2 evaluates the upper bound $\Psi(X;Y\|Z)$ for some settings. Finally, Section 4 discusses the tightness of the upper bound in the XOR case. Additionally, in Appendix A, we compute the secret key capacity for a class of sources where $Z$ is a function of $X$ and $P_{Y|X}$ corresponds to an erasure channel.
Notation and Definitions: Random variables are represented by uppercase letters, whereas their realizations are indicated by lowercase letters. It is assumed that all random variables possess finite alphabets. Sets are denoted by calligraphic letters. The notation $X-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-Z$ signifies that $X$ and $Z$ are conditionally independent given $Y$, which implies that $P_{XZ|Y}=P_{X|Y}P_{Z|Y}$. In this scenario, we assert that $X-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-Z$ constitutes a Markov chain. The sequences $(U_{1},U_{2},\ldots,U_{i})$ and $(U_{i},U_{i+1},\cdots,U_{n})$ are denoted as $U^{i}$ and $U_{i}^{n}$ respectively. For a sequence $X^{n}=(X_{1},X_{2},\cdots,X_{n})$, we use $X_{\backslash i}$ to denote $(X^{i-1},X_{i+1}^{n})$. The entropy of a random variable $X$ is represented as $H(X)$, while the mutual information between $X$ and $Y$ is denoted by $I(X;Y)$. We say $Q_{X}\ll P_{X}$ if $Q_{X}(x)=0$ whenever $P_{X}(x)=0$.
2 Main Results
The upper bound
$\hat{\Psi}(X;Y\|Z)$ $=\min_{J}I(X;Y|J)+I(X,Y;J|Z).$ (10)
satisfies
$\hat{\Psi}(X;Y\|Z)\leq\min(I(X;Y),I(X;Y|Z)).$ (11)
This follows from setting $J$ constant or $J=Z$. The main results of this section identify sufficient conditions for equality in (11). Applications will be provided in the next section.
Theorem 1.
Take some arbitrary $P_{X,Y,Z}$. Assume that one can find an auxiliary random variable $T$ on some arbitrary alphabet such that $I(T;Z|X,Y)=I(X;Y|T)=I(T;Z)=0$. Then
$\hat{\Psi}(X;Y\|Z)=I(X;Y).$
Proof.
Take an arbitrary $P_{J|X,Y,Z}$. Consider the coupling
$P_{T,X,Y,Z,J}\triangleq P_{T|X,Y,Z}P_{J|X,Y,Z}.$
Thus, we have
$I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.$
We have
$I(X;Y)-I(X;Y|J)$ $=I(J;X)-I(J;X|Y)$
$\leq I(J;T,X)-I(J;X|Y)$
$\leq I(J;T,X)+I(J;T,Y)-I(J;X,Y)$
=$\overset{(a)}{=}I(J;T,X)+I(J;T,Y)-I(J;X,Y,T)$
$=I(J;T,X)-I(J;X|T,Y)$
$\overset{(b)}{\leq}I(J;T)$
$\leq I(J,Z;T)$
$\overset{(c)}{\leq}I(J;T|Z)$
$\leq I(J;T,X,Y|Z)$
=$\overset{(d)}{=}I(J;X,Y|Z)$
where $(a)$ follows from $I(T;J|X,Y)=0$, $(b)$ follows from $I(X;Y|T)=0$, $(c)$ follows from $I(T;Z)=0$ and $(d)$ follows from $I(T;J|X,Y,Z)=0$. ∎
Remark 1.
$\hat{\Psi}(X;Y\|Z)=I(X;Y)$ is equivalent with the inequality
$I(X;Y|J)+I(X,Y;J|Z)\geq I(X;Y),\qquad\forall P_{J|X,Y,Z}.$
When $Z$ is a function of $(X,Y)$, this inequality can be equivalently written as
$I(J;X,Y,Z)\geq\frac{1}{2}I(J;X)+\frac{1}{2}I(J;Y)+\frac{1}{2}I(J;Z),\qquad\forall P_{J|X,Y,Z}.$
This condition holds if and only if $(1/2,1/2,1/2)$ belongs to the hypercontractivity ribbon for random variables $(X,Y,Z)$ (remember that the hypercontractivity ribbon for random variables $X_{1},X_{2},\cdots,X_{k}$ is defined as the set of $\lambda_{i}$ such that for any $P_{u|x_{[k]}}$ we have $I(X_{[k]};U)\geq\sum_{i}\lambda_{i}I(X_{i};U)$).
Remark 2.
Theorem 1 readily extends to quantum systems: take quantum systems $X,Y,Z,J$ with some arbitrary joint density $\rho_{X,Y,Z,J}$. The inequality
$I(X;Y|J)+I(X,Y;J|Z)\geq I(X;Y)$
holds if there is an extended state $\rho_{X,Y,Z,J,T}$ (whose reduced density on $X,Y,Z,J$ is the given state $\rho_{X,Y,Z,J}$) such that
$I(T;Z)=I(X;Y|T)=I(T;J|X,Y)=I(T;J|X,Y,Z)=0.$
The following lemma provides an equivalent condition for the assumption of Theorem 1 for the special case of $Z$ being a function of $(X,Y)$.
Lemma 1.
Consider the special case of $Z=f(X,Y)$ for some arbitrary function $f$. For a given joint distribution $P_{X,Y}$, there exists a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$ if and only if $P_{X,Y}$ lies in the convex hull of the set
$\mathcal{A}(P_{Z})=\left\{Q_{X,Y}:I_{Q}(X;Y)=0,~{}Q_{Z}=P_{Z}\right\}.$ (12)
Proof of this lemma can be found in Appendix B.
2.1 $Z$ is a function of $(X,Y)$
In this subsection, we assume that $Z$ is a function of $X$ and $Y$. The following lemma is immediate by the choice of $U=X,Y$ and $V=Z$ in (8) which would reduce the expression to that in (9).
Lemma 2.
We have $\hat{\Psi}(X;Y\|Z)={\Psi}(X;Y\|Z)$ if $Z$ is a function of $(X,Y)$.
We will explore the value of $\hat{\Psi}(X;Y\|Z)$ under various scenarios based on the alphabet sizes of $X$ and $Y$.
2.1.1 $X$ and $Y$ are binary
Assume that $X,Y\in\{0,1\}$ are binary random variables. Moreover,
$P_{X,Y}=\begin{bmatrix}p_{00}&p_{01}\\ p_{10}&p_{11}\end{bmatrix}.$
Let $X\oplus Y$ denote the XOR of $X$ and $Y$, and $X\wedge Y$ denote the AND of $X$ and $Y$. We consider the three functions $X+Y,X\oplus Y$ and $X\wedge Y$ as follows:
$Z_{1}=X+Y=\begin{cases}0&(X,Y)=(0,0)\\ 1&(X,Y)\in\{(0,1),(1,0)\}\\ 2&(X,Y)=(1,1)\end{cases}$
and
$Z_{2}=X\oplus Y=\begin{cases}1&(X,Y)\in\{(0,0),(1,1)\}\\ 0&(X,Y)\in\{(0,1),(1,0)\}\end{cases}$
and
$Z_{3}=X\wedge Y=\begin{cases}0&(X,Y)\in\{(0,1),(1,0),(0,0)\}\\ 1&(X,Y)=(1,1)\end{cases}.$
We have the following Theorem.
Theorem 2.
The following statements hold for binary random variables $X$ and $Y$:
  • For any $P_{X,Y}$, we have $\Psi(X;Y\|Z_{2})=I(X;Y).$
  • If $Cov(X,Y)\leq 0$, then $\Psi(X;Y\|Z_{1})=\Psi(X;Y\|Z_{3})=I(X;Y)$.
  • If $Cov(X,Y)>0$, then $\Psi(X;Y\|Z_{1})=\Psi(X;Y\|Z_{3})=0.$
Proof of this theorem can be found in Appendix B.
2.1.2 Ternary-Binary
In this section, we demonstrate the applicability of Theorem 1 beyond the binary case by extending the alphabet of $Y$ to $\mathcal{Y}=\{0,1,2\}$. We give a condition for $\Psi(X;Y\|Z)=I(X;Y)$ when $Z=(X+Y)~{}mod~{}2$.
Theorem 3.
Assume that $\mathcal{X}=\{0,1\},\mathcal{Y}=\{0,1,2\}$ and $\mathcal{Z}=\{0,1\}$, and
$Z=(X+Y)\mod 2=\begin{cases}0&(X,Y)\in\{(0,0),(1,1),(0,2)\}\\ 1&(X,Y)\in\{(0,1),(1,0),(1,2)\}\\ \end{cases}$ (13)
If
$\frac{P_{X,Y}(0,0)}{P_{Z}(0)}+\frac{P_{X,Y}(1,2)}{P_{Z}(1)}\leq 1$
and
$\frac{P_{X,Y}(0,2)}{P_{Z}(0)}+\frac{P_{X,Y}(1,0)}{P_{Z}(1)}\leq 1,$
then
$\Psi(X;Y\|Z)=I(X;Y).$
Proof of this theorem can be found in Appendix B.
Remark 3.
Note that when $P_{Y}(2)=0$ (i.e., $Y$ is binary), $Z$ in (13), will be the XOR of binary variables $X$ and $Y$. Therefore, Theorem 3 generalizes the first part of Theorem 2.
2.1.3 Arbitary Functions
This part considers random variables $X,Y$ and $Z$ with arbitrary alphabet sizes. We aim to determine all functions $f:\mathcal{X}\times\mathcal{Y}\mapsto\mathcal{Z}$ for which the equation $\Psi(X;Y|Z)=I(X;Y)$ holds for every joint distribution $P_{X,Y}$.
Theorem 4.
Take some fixed alphabets $\mathcal{X}$ and $\mathcal{Y}$. Fix some function $f:\mathcal{X}\times\mathcal{Y}\mapsto\mathcal{Z}$ and assume that $Z=f(X,Y)$. We have
$\Psi(X;Y\|Z)=I(X;Y),\qquad\forall P_{X,Y}$
if and only if either of the following holds:
  • $X$ and $Y$ are binary and we have the XOR mapping $f(0,0)=f(1,1)\neq f(1,0)=f(0,1)$,
  • $f(x,y)$ is a constant function.
Proof of Theorem 4 can be found in Appendix B.
Remark 4.
Theorem 4 shows that for every non-constant function other than the XOR function, there exists at least one distribution $P_{X,Y}$ such that $\Psi(X;Y\|Z)\neq I(X;Y)$.
3 On the gap between $S(X;Y\|Z)$, $\hat{\Psi}(X;Y\|Z)$ and $I(X;Y)$
Let us begin with the upper bound $\hat{\Psi}(X;Y\|Z)$ and $I(X;Y)$. It is clear that
$\hat{\Psi}(X;Y\|Z)\leq I(X;Y),\qquad\forall P_{X,Y,Z}.$
Assume that
$\hat{\Psi}(X;Y\|Z)=I(X;Y)$
for some $P_{X,Y,Z}$. Then, we claim that if $X^{\prime}=f_{1}(X)$, $Y^{\prime}=f_{2}(Y)$ and $Z^{\prime}=f_{3}(Z)$ for some functions $f_{1},f_{2},f_{3}$ such that $H(Z^{\prime}|X^{\prime},Y^{\prime})=0$, then, we also have
$\hat{\Psi}(X^{\prime};Y^{\prime}\|Z^{\prime})=I(X^{\prime};Y^{\prime}).$
We show this by establishing a more general result:
Theorem 5.
For any $P_{X,Y,Z,X^{\prime},Y^{\prime},Z^{\prime}}$, we have the following inequality:
$\hat{\Psi}(X;Y\|Z)-\hat{\Psi}(X^{\prime};Y^{\prime}|Z^{\prime})\leq I(XY;Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (14)
Remark 5.
Assume that $X^{\prime}=f_{1}(X)$, $Y^{\prime}=f_{2}(Y)$ and $Z^{\prime}=f_{3}(Z)$ for some functions $f_{1},f_{2},f_{3}$ such that $H(Z^{\prime}|X^{\prime},Y^{\prime})=0$. Then, (14) can be written as
$I(X;Y)-\hat{\Psi}(X,Y\|Z)\geq I(X^{\prime};Y^{\prime})-\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime}).$ (15)
If we further have $\hat{\Psi}(X,Y\|Z)=I(X;Y)$, we deduce that
$\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})\geq I(X^{\prime};Y^{\prime}).$ (16)
On the other hand, we always have
$\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})\leq I(X^{\prime};Y^{\prime}).$ (17)
Thus, $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})=I(X^{\prime};Y^{\prime})$.
Proof of Theorem 5 is given in Appendix B.
3.1 Theorem 5, a Conjecture and Lower bounds on $S(X;Y\|Z)$
It is not known whether there exists some $P_{X,Y,Z}$ such that $S(X;Y\|Z)\neq\hat{\Psi}(X;Y\|Z)$. If $S(X;Y\|Z)=\hat{\Psi}(X;Y\|Z)$ holds for all $P_{X,Y,Z}$, we must have the following for any $P_{XYZX^{\prime}Y^{\prime}Z^{\prime}}$:
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(XY;Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (18)
Observe that the above equation is valid when $X^{\prime}=X,Y^{\prime}=Y$ since from (6) we have
$S(X;Y\|Z)-S(X;Y\|Z^{\prime})\leq$ $S_{\text{ow}}(XY\rightarrow Z^{\prime}\|Z)\leq I(XY;Z^{\prime}|Z)$ (19)
We make the following conjecture:
Conjecture 1.
For any $P_{XYZX^{\prime}Y^{\prime}Z^{\prime}}$ we have
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq S_{\text{ow}}(XY\rightarrow Z^{\prime}\|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$
This conjecture generalizes (6) (it reduces to (6) by setting $X=X^{\prime}$ and $Y=Y^{\prime}$). If one can refute the conjecture, it will serve as evidence of $S(X;Y\|Z)\neq\hat{\Psi}(X;Y\|Z)$ for general distributions. If one can show the conjecture, one deduce the lower bound:
$S(X^{\prime};Y^{\prime}|Z^{\prime})\geq S(X;Y\|Z)-S_{\text{ow}}(XY\rightarrow Z^{\prime}|Z)-I(Y^{\prime}Z^{\prime};X|X^{\prime})-I(X^{\prime}Z^{\prime};Y|Y^{\prime})-I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$
This lower bound can be used to show that if
$S(X;Y\|Z)=I(X;Y)$
for some $P_{X,Y,Z}$ and if $X^{\prime}=f_{1}(X)$, $Y^{\prime}=f_{2}(Y)$ and $Z^{\prime}=f_{3}(Z)$ for some functions $f_{1},f_{2},f_{3}$ such that $H(Z^{\prime}|X^{\prime},Y^{\prime})=0$, then $S(X^{\prime};Y^{\prime}\|Z^{\prime})\geq I(X^{\prime};Y^{\prime})$ and hence,
$S(X^{\prime};Y^{\prime}\|Z^{\prime})=I(X^{\prime};Y^{\prime}).$
This shows that establishing the conjecture may require novel achievability (lower bound) ideas.
In Appendix C, we show that Conjecture 1 is equivalent to an apparently weaker form of the conjecture.
4 Is the upper bound $I(X;Y)$ tight for the XOR case?
Let $X$ and $Y$ be binary random variables with an arbitrary joint distribution, and set $Z=X\oplus Y$. We conclude from Theorem 2 that
$S(X;Y\|Z)\leq\Psi(X;Y\|Z)=\hat{\Psi}(X;Y\|Z)=I(X;Y).$
In other words, the upper bound $I(X;Y)$ cannot be improved using any existing upper bound. Note that $I(X;Y)$ is the maximum rate we could have achieved if Eve did not have access to the sequence $Z^{n}$ and could only monitor the public channel. We conjecture that the rate $I(X;Y)$ is no longer achievable (for a general $P_{X,Y}$) when Eve has access to the XOR of $X$ and $Y$.
Conjecture 2.
There exists some distribution $P_{X,Y}$ on binary random variables $X$ and $Y$ such that
$S(X;Y\|Z)<I(X;Y)$
where $Z=X\oplus Y$.
One idea to develop a better upper bound is to note that for every $n$ we have
$S(X;Y\|Z)=\frac{1}{n}S(X^{n};Y^{n}\|Z^{n})\leq\frac{1}{n}\Psi(X^{n};Y^{n}\|Z^{n})$
where $(X^{n},Y^{n},Z^{n})$ are $n$ i.i.d. repetitions of $(X,Y,Z)$. Therefore,
$S(X;Y\|Z)\leq\inf_{n}\left[\frac{1}{n}\Psi(X^{n};Y^{n}\|Z^{n})\right].$
The above inequality might potentially lead to better upper bounds because $\Psi(X^{n};Y^{n}\|Z^{n})$ allows for minimizing over all $P_{J|X^{n},Y^{n},Z^{n}}$ which is a larger space than that of the single-letter bound. The following proposition shows that at least when $Z$ is a function of $(X,Y)$, we do not obtain better upper bounds with this approach, i.e., the proposition shows that
$\inf_{n}\left[\frac{1}{n}\Psi(X^{n};Y^{n}\|Z^{n})\right]=\Psi(X;Y\|Z)$ (20)
as $\Psi(X;Y\|Z)=\hat{\Psi}(X;Y\|Z)$ when $Z$ is a function of $(X,Y)$ (however, we do not know if (20) holds when $Z$ is not a function of $(X,Y)$).
Proposition 2.
For every $n$, we have
$\hat{\Psi}(X;Y\|Z)=\frac{1}{n}\hat{\Psi}(X^{n};Y^{n}\|Z^{n})$
where $(X^{n},Y^{n},Z^{n})$ are $n$ i.i.d. repetitions of $(X,Y,Z)$.
Proof.
Take some arbitraray $P_{J|X^{n},Y^{n},Z^{n}}$ and let $J_{i}=(J,X^{i-1},Y^{i-1})$. Then, we have
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X^{n};Y^{n}|J)}+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X^{n}Y^{n};J|Z^{n})}$ $=\sum_{i}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i};Y_{i}|J_{i})+I(X^{i-1};Y_{i}|JY^{i-1})+I(Y^{i-1};X_{i}|JX^{i-1})}$
$\qquad+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i}Y_{i};J_{i}|Z_{i})+I(X_{i},Y_{i};Z_{\backslash i}|J_{i}Z_{i})}$
$\geq\sum_{i}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i};Y_{i}|J_{i})}+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(X_{i}Y_{i};J_{i}|Z_{i})}$
$\geq n\cdot\hat{\Psi}(X;Y\|Z),$ (21)
where the last equation follows from definition of $\hat{\Psi}(X;Y\|Z)$.
For the converse direction, assume that $J_{*}$ achieves the minimum of $I(X;Y|J)+I(X,Y;J|Z)$. Suppose $J$ consists of i.i.d copies of random variale $J_{*}$,i.e, $J=J_{*}^{n}$. We have
$I(X^{n};Y^{n}|J_{*}^{n})+I(X^{n}Y^{n};J_{*}^{n}|Z^{n})=n(I(X;Y|J_{*})+I(X,Y;J_{*}|Z))=n\hat{\Psi}(X;Y\|Z).$
Thus, $\hat{\Psi}(X^{n},Y^{n}\|Z^{n})\leq n\hat{\Psi}(X,Y\|Z)$. This observation and inequality (21) complete the proof. ∎
4.1 Controling $H(Z^{n}|F_{1:r})$
The following lemma shows that $H(Z^{n}|F_{1:r})$ cannot be equal to zero for any code that achieves a positive key rate:
Proposition 3.
Take some $P_{X,Y}(x,y)$ such that $P_{X,Y}(x,y)>0$ for all $x,y$. Assume that $H(Z^{n}|F_{1:r})=0$ where $F_{1:r}=(F_{1},F_{2},\cdots,F_{r})$ is the public communication. Then,
$H(X^{n},Y^{n}|F_{1:r})=0.$
We need the following definition:
Definition 1.
[16, 17] Given two pmfs $P_{X,Y}$ and $Q_{X,Y}$ on the alphabets $\mathcal{X}\times\mathcal{Y}$, the relation $Q_{X,Y}\preceq P_{X,Y}$ represents existence of some functions $a:\mathcal{X}\mapsto\mathbb{R}_{+}\cup\{0\}$ and $b:\mathcal{Y}\mapsto\mathbb{R}_{+}\cup\{0\}$ such that for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$ we have
$Q_{X,Y}(x,y)=a(x)b(y)P_{X,Y}(x,y)(x,y).$ (22)
Proof of Proposition 3.
Let us fix $F_{1:r}=f_{1:r}$. It is known that for any interactive communication, [^2] This property is similar to the rectangle property of communication complexity. we have
$P_{X^{n}Y^{n}|F_{1:r}=f_{1:r}}\preceq P_{X^{n}Y^{n}}.$
In other words,
$P(x^{n},y^{n}|F_{1:r}=f_{1:r})=P(x^{n},y^{n})a(x^{n})b(y^{n})$
for some functions $a(\cdot)$ and $b(\cdot)$.
Since $H(Z^{n}|F_{1:r}=f_{1:r})=0$, the value of $z^{n}$ must be fix when we fix $F_{1:r}=f_{1:r}$. This shows that for any $x_{1}^{n},y_{1}^{n}$ and $x_{2}^{n},y_{2}^{n}$ such that $x_{1}^{n}+y_{1}^{n}\neq x_{2}^{n}+y_{2}^{n}$, the probability of $p(x_{1}^{n},y_{1}^{n}|F_{1:r}=f_{1:r})$ and $p(x_{2}^{n},y_{2}^{n}|F_{1:r}=f_{1:r})$ cannot be positive at the same time. Thus,
$0=P(x_{1}^{n},y_{1}^{n}|F_{1:r}=f_{1:r})P(x_{2}^{n},y_{2}^{n}|F_{1:r}=f_{1:r})=P(x_{1}^{n},y_{1}^{n})P(x_{2}^{n},y_{2}^{n})a(x_{1}^{n})b(y_{1}^{n})a(x_{2}^{n})b(y_{2}^{n}).$
Since $P_{X,Y}(x,y)>0$ for all $x,y$, we must have $a(x_{1}^{n})b(y_{1}^{n})a(x_{2}^{n})b(y_{2}^{n})=0$.
Let $y^{n}$ be such that $b(y^{n})>0$. Take two distinct sequences $x_{1}^{n},x_{2}^{n}$ and set $y_{1}^{n}=y_{2}^{n}=y^{n}$. This choice is valid since $x_{1}^{n}+y_{1}^{n}\neq x_{2}^{n}+y_{2}^{n}$. Thus, we get
$a(x_{1}^{n})a(x_{2}^{n})b(y^{n})=0.$
Since we assumed $b(y^{n})>0$, we get that either $a(x_{1}^{n})$ or $a(x_{2}^{n})$ must be zero. Therefore, $a(x^{n})=0$ for all but one sequence $x^{n}$, and similarly we can show that $b(y^{n})=0$ for all but one sequence $y^{n}$. Thus,
$P(x^{n},y^{n}|F_{1:r}=f_{1:r})=P(x^{n},y^{n})a(x^{n})b(y^{n})$
is positive for only one pair $(x^{n},y^{n})$. Thus,
$H(X^{n},Y^{n}|F_{1:r}=f_{1:r})=0.$
Proposition 3 motivates the following definition:
Definition 2.
$S_{\Delta}(X;Y\|Z)$ is supremum of secret key rate corresponding to protocols satisfying
$\lim_{n\rightarrow\infty}\frac{1}{n}H(Z^{n}|F_{1},F_{2},\cdots,F_{r})\geq\Delta.$
The non-interactive capacity $S_{\text{ni},\Delta}(X;Y\|Z)$ is defined similarly.
Remark 6.
Observe that $\Delta\mapsto S_{\Delta}(X;Y\|Z)$ is a non-increasing function. Moreover,
$S(X;Y\|Z)=\lim_{\Delta\rightarrow 0}S_{\Delta}(X;Y\|Z).$
Our strategy for improving the upper bound on $S(X;Y\|Z)$ is as follows:
  1. 1.
    Compute an upper bound for $S_{\Delta}(X;Y\|Z)$ for any given $\Delta>0$. We will provide such an upper bound in Theorem 6 below.
  2. 2.
    Show that $S(X;Y\|Z)=S_{\Delta^{*}}(X;Y\|Z)$ for some $\Delta^{*}>0$. Intuitively we expect this to be true because of Proposition 3. We only show this part for the non-interactive setup for the XOR example.
  3. 3.
    Use the upper bound on $S_{\Delta^{*}}(X;Y\|Z)$ to obtain an upper bound on $S(X;Y\|Z)$. We are unable to perform this last step for the non-interactive setting because, even though the upper bound on $S_{\Delta^{*}}(X;Y\|Z)$ is theoretically computable, unfortunately, the number of free variables is large and it is not easy to obtain reliable numerical results.
We begin with task 1:
Theorem 6.
$S_{\Delta}(X;Y\|Z)$ has an upper bound as follows:
$S_{\Delta}(X;Y\|Z)\leq\Psi_{\Delta}(X;Y\|Z)\triangleq\inf_{J}\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(X;Y|J)+I(U;J|V)-I(U;Z|V),$ (23)
where $P_{X,Y,Z,J,U,V}$ satisfies the following constraint
$I(V;Z)\leq I(V;J)+H(Z)-\Delta,$ (24)
where in the inner maximum $U$ and $V$ satisfy the cardinality bounds $|\mathcal{V}|\leq|\mathcal{X}||\mathcal{Y}|+1$ and $|\mathcal{U}|\leq|\mathcal{X}||\mathcal{Y}|$.
Moreover, if $Z=f(X,Y)$, in computing the upper bound we can assume that $H(X,Y|U,V,Z)=0$.
Remark 7.
Observe that $\Delta\mapsto\Psi_{\Delta}(X;Y\|Z)$ is a non-increasing function. When $\Delta=0$, $\Psi_{\Delta}(X;Y\|Z)$ reduces to $\Psi(X;Y\|Z)$ as the only difference between the optimization problems in (7) and (23) is the additional constraint defined in (24). This constraint holds trivially when $\Delta=0$. This constraint imposes a limitation on the distribution $P_{U,V|X,Y}$.
As observed earlier in computing $\Psi(X;Y\|Z)$ when $Z$ is a function of $(X,Y)$, an optimizer for
$\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(U;J|V)-I(U;Z|V)$
is $U=X,Y$ and $V=Z$, which yields the value $I(X,Y;J|Z)$. However, for a positive $\Delta$, this choice of $(U,V)$ may not be feasible in (24).
Proof of Theorem 6 is given in Appendix B.
We now turn to the second task. The following theorem provides a lower bound on $H(Z^{n}|F_{1},F_{2})$ for a capacity achieving code. Some of the manipulations inside the proof are similar to the one given in [18] in the context of a different distributed source coding problem.
Theorem 7.
Suppose that $X$ and $Y$ are binary random variables, and let $Z=X\oplus Y$. Let
$\Delta_{1}$ $=\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}H(Z|U)-H(X|U),~{}\Delta_{2}=\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-(X,Z)}H(Z|V)-H(X|V),$ (25)
$\bar{\Delta}$ $=\max\{\Delta_{1},\Delta_{2}\}.$ (26)
Then, every non-interactive code with public messages $F_{1}$ and $F_{2}$ satisfies
$\frac{H(Z^{n}|F_{1},F_{2})}{n}\geq\frac{\bar{\Delta}}{2}+\frac{I(K_{A};K_{B}|F_{1},F_{2})}{2n}.$ (27)
Corollary 1.
Let
$\Delta=\frac{\bar{\Delta}+S_{\text{ni}}(X;Y\|Z)}{2}.$
Then, we can conclude:
$S_{\text{ni}}(X;Y\|Z)=S_{\text{ni},\Delta}(X;Y\|Z).$ (28)
Proof.
We have
$I(K_{A};K_{B}|F_{1},F_{2})$ $\leq I(X^{n};Y^{n}|F_{1},F_{2})$
$\leq H(X^{n}Y^{n}|F_{1},F_{2})$
$=nH(X,Y)-I(F_{1},F_{2};X^{n},Y^{n})$
This yields
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})\leq H(X,Y)-\frac{1}{n}I(K_{A};K_{B}|F_{1},F_{2}).$
To prove the desired inequality, it suffices to show the following inequalities:
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2H(Z^{n}|F_{1},F_{2})}{n}$ $\geq H(X,Y)+\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}H(Z|U)-H(Y|U),$ (29)
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2H(Z^{n}|F_{1},F_{2})}{n}$ $\geq H(X,Y)+\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-(X,Z)}H(Z|V)-H(X|U),$ (30)
as combining inequalities (29) and (30) gives:
$\frac{1}{n}I(F_{1},F_{2};X^{n},Y^{n})+\frac{2H(Z^{n}|F_{1},F_{2})}{n}\geq H(X,Y)+\bar{\Delta}.$ (31)
We only prove inequality (29), as the proof of (30) is similar. Let $U_{i}=(F_{1},Y^{i-1},Z_{i+1}^{n})$, $V_{i}=(F_{2},X^{i-1},Z_{i+1}^{n})$ be new auxiliary random variables. We have:
$I(F_{1},F_{2};X^{n},Y^{n})+2H(Z^{n}|F_{1},F_{2})$ $=I(F_{1};X^{n},Y^{n})+I(F_{2};Y^{n},X^{n}|F_{1})+2H(Z^{n}|F_{1},F_{2})$
$=I(F_{1};X^{n})+I(F_{2};Y^{n},Z^{n}|F_{1})+2H(Z^{n}|F_{1},F_{2})$
$=I(F_{1};X^{n})+I(F_{2};Z^{n}|F_{1})+I(F_{2};Y^{n}|F_{1},Z^{n})+2H(Z^{n}|F_{1},F_{2})$
$=I(F_{1};X^{n})+H(Z^{n}|F_{1})+I(F_{2};Y^{n}|F_{1},Z^{n})+H(Z^{n}|F_{1},F_{2})$
=$\overset{(a)}{=}I(F_{1};X^{n})+H(Y^{n}|F_{1})+H(Z^{n}|F_{1},Y^{n})+I(F_{2};Y^{n}|F_{1},Z^{n})$
$\qquad+H(Z^{n}|F_{1})-H(Y^{n}|F_{1})+I(Y^{n};Z^{n}|F_{1},F_{2})$
=$\overset{(b)}{=}H(X^{n},Y^{n})+I(F_{2};Y^{n}|F_{1},Z^{n})+H(Z^{n}|F_{1})-H(Y^{n}|F_{1})$
$\qquad+I(Z^{n};Y^{n}|F_{1},F_{2})$
$\overset{(c)}{\geq}nH(X,Y)+H(Z^{n}|F_{1})-H(Y^{n}|F_{1})$
=$\overset{(d)}{=}nH(X,Y)+\sum_{i}H(Z_{i}|F_{1},Y^{i-1},Z_{i+1}^{n})-H(Y_{i}|F_{1},Y^{i-1},Z_{i+1}^{n})$
$=nH(X,Y)+\sum_{i}H(Z_{i}|U_{i})-H(Y_{i}|U_{i}),$ (32)
where $(a)$ follows from $H(Z^{n}|F_{1},F_{2},Y^{n})=H(Z^{n}|F_{1},Y^{n})$, which is equivalent with $H(X^{n}|F_{1},F_{2},Y^{n})=H(X^{n}|F_{1},Y^{n})$ which holds due to the Markov chain $F_{2},Y^{n}-\!\!\!\!\minuso\!\!\!\!-X^{n}-\!\!\!\!\minuso\!\!\!\!-F_{1}$, and for $(b)$ we use the following
$I(F_{1};X^{n})+H(Y^{n}|F_{1})+H(Z^{n}|F_{1},Y^{n})$ $=I(F_{1};X^{n})+H(Y^{n}|F_{1})+H(X^{n}|F_{1},Y^{n})$
$=I(F_{1};X^{n})+H(X^{n},Y^{n}|F_{1})$
$=H(X^{n},Y^{n}).$
Also, in the above, $(c)$ follows from the non-negativity of conditional mutual information, and $(d)$ follows from the Marton-Körner identity.
Let $Q$ be the uniform distribution over the set $\{1,2,\cdots,n\}$, independent of $X,Y$ and $F_{1:r}$, and let $X=X_{Q},Y=Y_{Q}$ and $U=(U,Q)$. It is clear that we have the Markov chain $U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-Y$. Thus, inequality (32) simplifies to the desired bound. ∎
We note that the optimization problems $\min_{U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-Y}H(Z|U)-H(Y|U)$ and $\min_{V-\!\!\!\!\minuso\!\!\!\!-Y-\!\!\!\!\minuso\!\!\!\!-X}H(Z|V)-H(X|V)$ also arise in many places, including the “modulo 2 sum” problem in [18]. For instance, the following is known:
Example 1 (Proposition $1$ and $2$ of [18]).
Let $P_{X}(0)=x,P_{X|Y}(0|0)=c$ and $P_{X|Y}(1|1)=d$. The value of $\bar{\Delta}$ is 00 if $(1-2c)(1-2d)\leq 0$. Additionally, by setting $x=\frac{\sqrt{d(1-d)}}{\sqrt{d(1-d)}+\sqrt{c(1-c)}}$, the value of $\bar{\Delta}$ is $2H(Z)-H(X,Y)$ if either $c=d$ or $(1-2c)(1-2d)>0$ and $c\neq d$.
5 Acknowgement
Prior to the first author’s involvement in this problem, the second author designated this project as the undergraduate thesis topic for Yicheng Cui. During the initial stages of this research, Cui made some progress on the XOR and AND patterns, employing different (and weaker) techniques than those presented in this work.
The authors would like to also thank Mr. Chin Wa Lau for his help in proving Proposition 2 and for helping Yicheng Cui in his undergraduate thesis.
Appendix A Capacity of a class of source distributions
Assume that $Z=g(X)$ and the channel from $X$ to $Y$ is erasure, i.e., $Y=X$ with probability $1-\epsilon$ and $\{e\}$ otherwise. Note that random variables $X,Y$ and $Z$ do not always form any Markov chain in any order, and the key capacity for this class is not known in the literature.
We claim that $S(X;Y\|Z)=I(X;Y|Z)=(1-\epsilon)H(X|Z)$ is achievable by one-way communication from Alice to Bob. Note that
$S_{\text{ow}}(X\rightarrow Y\|Z)=\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}I(U;Y|V)-I(U;Z|V).$ (33)
Consider $V$ constant, and by the functional representation lemma, let $U$ be such that $I(Z;U)=0$ and $H(X|Z,U)=0$. Then, note that $I(U;Z|Y)=0$ since when $Y$ is $X$ it holds and when $Y$ is $e$, it also holds. Then, we have
$I(U;Y|V)-I(U;Z|V)$ $=I(U;Y)-I(U;Z)$ (34)
$=I(U;Y)$ (35)
$=I(U;YZ)$ (36)
$=I(U;Y|Z)$ (37)
$=I(X;Y|Z)-I(X;Y|Z,U)$ (38)
$=I(X;Y|Z).$ (39)
This completes the proof.
Appendix B Some of the proofs
B.1 Proof of Lemma 1
Proof of Lemma 1.
Note that $I(Z;T|X,Y)=0$ holds trivially when $Z$ is a function of $(X,Y)$. Assume that for given $P_{X,Y}$ there exists a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$. Define $Q_{X,Y}^{(t)}=P(X,Y|T=t)$. From $I(X;Y|T)=0$, it follows that $I_{Q^{(t)}}(X;Y)=0$. Additionally, for any $t\in\mathcal{T}$ and $z\in\mathcal{Z}$, we have
$Q^{(t)}(Z=z)=\sum_{x,y\in\mathcal{R}_{z}}Q^{(t)}(X=x,Y=y)=\sum_{x,y\in\mathcal{R}_{z}}P(X=x,Y=y|T=t)=P(Z=z|T=t)\overset{(a)}{=}P(Z=z),$
where $\mathcal{R}_{z}=\left\{(x,y)\in\mathcal{X}\times\mathcal{Y}:z=f(x,y)\right\}$ and $(a)$ follows from $I(Z;T)=0$. Hence, we conclude that $P_{X,Y}\in conv(A)$.
For the converse direction, assume that $P_{X,Y}\in conv(A)$. Therefore, we can represent $P_{X,Y}$ as a convex combination of elements from $A$:
$P_{X,Y}=\sum_{t}\omega_{t}Q_{X,Y}^{(t)},~{}Q_{X,Y}^{(t)}\in A,$ (40)
where $\omega_{t}\geq 0$ and $\sum_{t}\omega_{t}=1$. Define $P(X,Y|T=t)=Q_{X,Y}^{(t)}$. Since $Q^{(t)}\in A$, it follows that $I_{P}(X;Y)=I_{Q^{(t)}}(X;Y)=0$. Next,
$P(Z=z|T=t)=\sum_{x,y\in\mathcal{R}_{z}}P(X=x,Y=y|T=t)=\sum_{x,y\in\mathcal{R}_{z}}Q^{(t)}(X=x,Y=y)=Q(Z=z)\overset{(a)}{=}P(Z=z)$
where $(a)$ holds since $Q^{(t)}\in A$. Thus form above we have $I_{P}(Z;T)=0$. This completes the proof. ∎
B.2 Proof of Theorem 2
Proof of Theorem 2.
The condition $Cov(X,Y)\leq 0$ is equivalent with $p_{00}p_{11}\leq p_{01}p_{10}$.
For the first part, since
$\Psi(X;Y\|Z_{3})\geq\Psi(X;Y\|Z_{1})$
it suffices to show that if $p_{00}p_{11}\leq p_{01}p_{10}$, then $\Psi(X;Y\|Z_{1})=I(X;Y)$ and if $p_{00}p_{11}>p_{01}p_{10}$, then $\Psi(X;Y\|Z_{3})=0.$
Proof of $\Psi(X;Y\|Z_{1})=I(X;Y)$ if $p_{00}p_{11}\leq p_{01}p_{10}$: The condition $p_{00}p_{11}\leq p_{01}p_{10}$ implies that the polynomial
$x^{2}-(p_{01}+p_{10})x+p_{00}p_{11}=0$
has two non-negative real roots. Let $x_{0}\leq x_{1}$ denote the roots. Note that $x_{0}x_{1}=p_{00}p_{11}$ and $x_{0}+x_{1}=p_{01}+p_{10}$. One can verify that $p_{01}$ belongs to $[x_{0},x_{1}]$. Then, one can find $\lambda\in[0,1]$ such that
$p_{01}=\lambda x_{0}+(1-\lambda)x_{1}.$
since $x_{0}+x_{1}=p_{01}+p_{10}$, we obtain
$p_{10}=\lambda x_{1}+(1-\lambda)x_{0}.$
Since $\Psi(X;Y\|Z_{1})=\hat{\Psi}(X;Y\|Z_{1})$, to show that $\hat{\Psi}(X;Y\|Z_{1})=I(X;Y)$, we use Theorem 1. It suffices to identify a $T$ such that $I(T;Z_{1}|X,Y)=I(T;Z_{1})=I(X;Y|T)=0$. The condition $I(T;Z_{1}|X,Y)=0$ holds for any $P_{T|X,Y}$ because $Z_{1}$ is a function of $(X,Y)$. Let $T$ be a binary random variable satisfying $P(T=1)=\lambda$ and $P(T=0)=1-\lambda$ and
$P(X=1,Y=1|T=t)=p_{11},\qquad t\in\{0,1\}$
$P(X=0,Y=0|T=t)=p_{00},\qquad t\in\{0,1\}$
$P(X=0,Y=1|T=0)=P(X=1,Y=0|T=1)=x_{1}.$
$P(X=0,Y=1|T=1)=P(X=1,Y=0|T=0)=x_{0}$
This yields the desired result.
Proof of $\Psi(X;Y\|Z_{3})=0$ if $p_{00}p_{11}>p_{01}p_{10}$: we identify a binary random variable $J^{*}$ such that $I(X;Y|J^{*})=I(X,Y;J^{*}|Z_{3})=0$. This would imply that
$\Psi(X;Y\|Z_{3})\leq I(X;Y|J^{*})+I(X,Y;J^{*}|Z_{3})=0.$
Let
$\alpha=\dfrac{p_{00}}{p_{00}-(p_{00}p_{11}-p_{01}p_{10})}.$
Under condition $p_{00}p_{11}-p_{01}p_{10}>0$ we have $\alpha\geq 1$. Suppose that
$P_{X,Y|J^{*}=0}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$
and
$P_{X,Y|J^{*}=1}=\begin{bmatrix}\alpha p_{00}&\alpha p_{01}\\ \alpha p_{10}&1-\alpha(1-p_{11})\end{bmatrix}$
and $P(J^{*}=1)=\frac{1}{\alpha}$. One can directly verify that
$P_{X,Y}=P(J^{*}=1)P_{X,Y|J^{*}=1}+P(J^{*}=0)P_{X,Y|J^{*}=0}.$
Moreover, for this choice of $J^{*}$, we have $I(X;Y|J^{*})=I(X,Y;J^{*}|Z)=0$.
For the second part, we consider two cases: if $p_{00}p_{11}\leq p_{01}p_{10}$, we use the fact that $\Psi(X;Y\|Z_{1})=I(X;Y)$ and
$\Psi(X;Y\|Z_{2})\geq\Psi(X;Y\|Z_{1})=I(X;Y).$
In the case of $p_{00}p_{11}\leq p_{01}p_{10}$, consider the following random variable:
$\hat{Z}_{1}=\begin{cases}0&(X,Y)=(0,1)\\ 1&(X,Y)=(1,0)\\ 2&(X,Y)\in\{(0,0),(1,1)\}\end{cases}$
Then, by symmetry, from the first part we deduce that in the case of $p_{00}p_{11}\leq p_{01}p_{10}$, we have $\Psi(X;Y\|\hat{Z}_{1})=I(X;Y)$. Since
$\Psi(X;Y\|Z_{2})\geq\Psi(X;Y\|\hat{Z}_{1})=I(X;Y)$
we get the desired result. ∎
B.3 Proof of Theorem 3
Proof of Theorem 3.
Note that when $P(Y=2)=0$, i.e., when $Y\in\{0,1\}$, the random variable $Z$ can be expressed as $Z=X\oplus Y$. In this case, proof of Theorem 2 shows that one can find a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$. Lemma 1 would then imply any $P_{X,Y}$ satisfying $P_{Y}(2)=0$ belongs to the convex hull of the set $\mathcal{A}(P_{Z})$ as defined in (12). More specifically, let $\alpha=P_{Z}(0)$ and $\bar{\alpha}=P_{Z}(1)=1-\alpha$ and let us denote $\mathcal{A}(P_{Z})$ by $\mathcal{A}(\alpha)$. Then, for any $r,t\in[0,1]$, the following joint distribution belongs to the convex hull of the set $\mathcal{A}(\alpha)$:
[…]
Observe that when $P(Y=0)=0$, the mapping from $X$ and $Y\in\{1,2\}$ to $Z$ is again an XOR map. A similar argument shows that then imply any $P_{X,Y}$ satisfying $P_{Y}(0)=0$ belongs to the convex hull of the set $\mathcal{A}(\alpha)$. In other words, for any $s,\ell\in[0,1]$ the following joint distribution belongs to the convex hull of the set $\mathcal{A}(\alpha)$:
[…]
Observe that for any $\omega\in[0,1]$ the linear combination
$\omega B_{r,t}(x,y)+\bar{\omega}C_{s,\ell}(x,y)$
must also belong to the convex hull of the set $\mathcal{A}(\alpha)$.
Let $p_{ij}=P_{X,Y}(i,j)$. We have $\alpha=p_{00}+p_{02}+p_{11}$. Set
$r=\min\left\{1,\frac{\bar{\alpha}p_{00}}{\alpha p_{10}}\right\}$
$\omega=\frac{p_{00}}{\alpha r},~{}t=\frac{\alpha p_{10}}{\bar{\alpha}p_{00}}r,~{}s=\frac{\alpha p_{12}r}{\bar{\alpha}(\alpha r-p_{00})},~{}\ell=\frac{p_{02}r}{\alpha r-p_{00}}.$
One can then verify that
$P_{X,Y}(x,y)=\omega B_{r,t}(x,y)+\bar{\omega}C_{s,\ell}(x,y)$
belongs to the convex hull of the set $\mathcal{A}(\alpha)$. The conditions
$\frac{p_{00}}{\alpha}+\frac{p_{12}}{\bar{\alpha}}\leq 1,~{}\frac{p_{02}}{\alpha}+\frac{p_{10}}{\bar{\alpha}}\leq 1.$
imply that
$\omega,t,r,s,\ell\in[0,1].$
Consequently, one can identify a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$ holds for $P_{X,Y}$ provided that the conditions
$\frac{p_{00}}{\alpha}+\frac{p_{12}}{\bar{\alpha}}\leq 1,~{}\frac{p_{02}}{\alpha}+\frac{p_{10}}{\bar{\alpha}}\leq 1$
are satisfied.
B.4 Proof of Theorem 4
Proof of Theorem 4.
Let the alphabets of $X$ and $Y$ take values in $\{0,1,2,\cdots,s_{1}-1\}$ and $\{0,1,2,\cdots,s_{2}-1\}$ respectively. Let $Z=f(X,Y)$. The function $f$ corresponds to a table
A table is considered valid if the equality $\Psi(X;Y|Z)=I(X;Y)$ holds for all distributions $P_{X,Y}$. Since any distribution over a subset of $\mathcal{X}\times\mathcal{Y}$ can be extended to a distribution over $\mathcal{X}\times\mathcal{Y}$, it follows that if a table is valid, all of its sub-tables will also be valid. Consider a $2\times 2$ table. As stated in Theorem 2, the equality $\Psi(X;Y|Z)=I(X;Y)$ does not hold for any arbitrary distribution $P_{X,Y}$ when the functions are AND, OR, or $Z_{1}$. It is also evident that if $Z=f(X,Y)$ is a one-to-one map, $\Psi(X;Y|Z)\leq I(X;Y|Z)=0$ and the equality $\Psi(X;Y|Z)=I(X;Y)$ would not hold for every $P_{X,Y}$. Consequently, for $2\times 2$ tables, the following combinations are invalid if $A,B,C$ and $D$ are distinct symbols:
and the valid $2\times 2$ tables are limited to the following:
[…]
Now, consider a general table $T$ of arbitrary size $s_{1}\times s_{2}$ where $(s_{1},s_{2})\neq(2,2)$. Since every $2\times 2$ sub-table must be valid, we conclude that for every $i_{1}\neq i_{2}$ and $j_{1}\neq j_{2}$, we must either have
$T(i_{1},j_{1})=T(i_{1},j_{2})=T(i_{2},j_{1})=T(i_{2},j_{2})$
or
$T(i_{1},j_{1})=T(i_{2},j_{2})\neq T(i_{2},j_{1})=T(i_{1},j_{2})$
In both cases $T(i_{1},j_{1})=T(i_{2},j_{2})$ for every $i_{1}\neq i_{2}$ and $j_{1}\neq j_{2}$. In other words, if we move from one cell to another cell by changing both the row and the column, the value of the new cell must be equal to the value of the old cell. Note that in a $s_{1}\times s_{2}$ table where $(s_{1},s_{2})\neq(2,2)$, we can move from a cell to all the other cells by a sequence of such moves. For instance, in a $2\times 3$ table, we can move from $(1,1)$ to $(1,2)$ by the following sequence:
$(1,1)\rightarrow(2,2)\rightarrow(3,1)\rightarrow(1,2).$
This shows that when $(s_{1},s_{2})\neq(2,2)$, $T(i_{1},j_{1})=T(i_{2},j_{2})$ for all $i_{1},i_{2},j_{1},j_{2}$. ∎
B.5 Proof of Theorem 5
Proof of Theorem 5.
Let $P_{J|X^{\prime}Y^{\prime}Z^{\prime}}$ be a minimizer for the optimization problem in $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})$, i.e.,
$\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})=I(X^{\prime};Y^{\prime}|J)+I(X^{\prime},Y^{\prime};J|Z^{\prime}).$
We know that such a minimizer exists because of the cardinality bound on $J$ is available for $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})$.
We define the joint distribution between $J$ and $(X,Y,Z,X^{\prime},Y^{\prime},Z^{\prime})$ as $P_{J|X^{\prime}Y^{\prime}Z^{\prime}}P_{X^{\prime}Y^{\prime}Z^{\prime}XYZ}$. This construction implies the Markov chain
$(X,Y,Z)-\!\!\!\!\minuso\!\!\!\!-(X^{\prime},Y^{\prime},Z^{\prime})-\!\!\!\!\minuso\!\!\!\!-J.$ (75)
Since
$\hat{\Psi}(X;Y\|Z)\leq I(X;Y|J)+I(X,Y;J|Z)$
it suffices to show that
$I(X;Y|J)+I(X,Y;J|Z)-\hat{\Psi}(X^{\prime};Y^{\prime}\|Z^{\prime})$ $=I(X;Y|J)+I(X,Y;J|Z)-I(X^{\prime};Y^{\prime}|J)-I(X^{\prime},Y^{\prime};J|Z^{\prime})$
$\leq I(XY;Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (76)
This inequality follows from adding the following three inequalities:
$I(X;Y|J)+I(XY;J|Z)-I(XX^{\prime};YY^{\prime}|J)-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$ $\leq I(XY;Z^{\prime}|Z),$ (77)
$I(XX^{\prime};YY^{\prime}|J)+I(XX^{\prime}YY^{\prime};J|Z^{\prime})-I(XX^{\prime};Y^{\prime}|J)-I(XX^{\prime}Y^{\prime};J|Z^{\prime})$ $\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime}),$ (78)
$I(XX^{\prime};Y^{\prime}|J)+I(XX^{\prime}Y^{\prime};J|Z^{\prime})-I(X^{\prime};Y^{\prime}|J)-I(X^{\prime}Y^{\prime};J|Z^{\prime})$ $\leq I(Y^{\prime}Z^{\prime};X|X^{\prime}).$ (79)
The inequality (77) follows from
$I(X;Y|J)+I(XY;J|Z)-I(XX^{\prime};YY^{\prime}|J)-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$
$\leq I(XY;JZ^{\prime}|Z)-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$
$\leq I(XY;Z^{\prime}|Z)+I(XYZ;J|Z^{\prime})-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$
$\leq I(XY;Z^{\prime}|Z)$
where the last step follows from (75).
The inequality (78) follows from
$I(XX^{\prime};YY^{\prime}|J)+I(XX^{\prime}YY^{\prime};J|Z^{\prime})-I(XX^{\prime};Y^{\prime}|J)-I(XX^{\prime}Y^{\prime};J|Z^{\prime})$
$=I(XX^{\prime};Y|JY^{\prime})+I(Y;J|XX^{\prime}Y^{\prime}Z^{\prime})$
$=I(XX^{\prime};Y|JY^{\prime})-I(Y;J|XX^{\prime}Y^{\prime}Z^{\prime})$ (80)
$\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime})-I(J;Y|Y^{\prime})$
$\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime})$
where (80) follows from (75).
Finally, the inequality (79) similarly follows from
$I(XX^{\prime};Y^{\prime}|J)+I(XX^{\prime}Y^{\prime};J|Z^{\prime})-I(X^{\prime};Y^{\prime}|J)-I(X^{\prime}Y^{\prime};J|Z^{\prime})$ (81)
$=I(X;Y^{\prime}|X^{\prime}J)+I(X;J|X^{\prime}Y^{\prime}Z^{\prime})$ (82)
$=I(X;Y^{\prime}|X^{\prime}J)-I(X;J|X^{\prime}Y^{\prime}Z^{\prime})$ (83)
$\leq I(Y^{\prime}Z^{\prime};X|X^{\prime})-I(J;X|X^{\prime})$ (84)
$\leq I(Y^{\prime}Z^{\prime};X|X^{\prime})$ (85)
This completes the proof. ∎
B.6 Proof of Theorem 6
Proof of Theorem 6.
Let $F_{1:r}=(F_{1},\cdots,F_{r})$. Take some arbitrary $P_{J|X,Y,Z}$ and suppose $(X^{n},Y^{n},Z^{n},J^{n})$ are $n$ i.i.d. copies according to $P_{X,Y,Z}P_{J|X,Y,Z}$. Define the auxiliary random variables as $V_{i}=(J^{i-1},Z_{i+1}^{n},F_{1:r})$ and $U_{i}=K_{A}$. The following chains of equations follows:
$H(K_{A})$ $=I(K_{A};Z^{n},F_{1:r})+H(K_{A}|Z^{n},F_{1:r})$ (86)
$\overset{(a)}{\leq}n\epsilon_{n}+H(K_{A}|Z^{n},F_{1:r})$
$=n\epsilon_{n}+H(K_{A}|Z^{n},F_{1:r})-H(K_{A}|J^{n},F_{1:r})+H(K_{A}|J^{n},F_{1:r})$
=$\overset{(b)}{=}n\epsilon_{n}+\sum_{i=1}^{n}I(K_{A};J_{i}|J^{i-1},Z_{i+1}^{n},F)-I(K_{A};Z_{i}|J^{i-1},Z_{i+1}^{n},F)+H(K_{A}|J^{n},F_{1:r})$
=$\overset{(c)}{=}n\epsilon_{n}+n\delta_{n}+\sum_{i=1}^{n}I(U_{i};J_{i}|V_{i})-I(U_{i};Z_{i}|V_{i})+I(K_{A};K_{B}|J^{n},F_{1:r})$
$\overset{(d)}{\leq}n\epsilon_{n}+n\delta_{n}+\sum_{i=1}^{n}I(U_{i};J_{i}|V_{i})-I(U_{i};Z_{i}|V_{i})+I(X^{n};Y^{n}|J^{n},F_{1:r})$
$\overset{(e)}{\leq}n\epsilon_{n}+n\delta_{n}+\sum_{i=1}^{n}I(U_{i};J_{i}|V_{i})-I(U_{i};Z_{i}|V_{i})+I(X^{n};Y^{n}|J^{n})$
$\leq n\epsilon_{n}+n\delta_{n}+\sum_{i=1}^{n}I(U_{i};J_{i}|V_{i})-I(U_{i};Z_{i}|V_{i})+nI(X;Y|J)$ (87)
where $\epsilon_{n},\delta_{n}\rightarrow 0$ as $n\rightarrow\infty$; steps (a) and (c) are derived using Fano’s inequality, while (b) follows from the Marton-Körner identity, and (d) is a consequence of the Data Processing Inequality (DPI). To establish (e), we proceed as follows:
$I(X^{n};Y^{n}|J^{n})$ =$\overset{(a)}{=}I(X^{n},F_{1};Y^{n}|J^{n})\overset{(b)}{\geq}I(X^{n};Y^{n}|J^{n},F_{1})\overset{(c)}{=}I(X^{n};Y^{n},F_{2}|J^{n},F_{1})\overset{(d)}{\geq}I(X^{n};Y^{n}|J^{n},F_{1},F_{2})$
$=I(X^{n},F_{3};Y^{n}|J^{n},F_{1},F_{2})\geq I(X^{n};Y^{n}|J^{n},F_{1},F_{2},F_{3})=\cdots\geq I(X^{n};Y^{n}|J^{n},F_{1:r}),$
where $(a)$ and $(b)$ follow from Markov chains $(Y^{n},F^{i-1})-\!\!\!\!\minuso\!\!\!\!-X^{n}-\!\!\!\!\minuso\!\!\!\!-F_{i}$ for odd $i$ and $(X^{n},F^{i-1})-\!\!\!\!\minuso\!\!\!\!-Y^{n}-\!\!\!\!\minuso\!\!\!\!-F_{i}$ for even $i$ respectively. Also $(b)$ and $(d)$ are derived from chain rule.
Let $Q$ be uniformly distributed over $\{1,2,\cdots,n\}$ and independent of $X,Y,Z,J,F_{1:r}$. Define the random variables $V=(V_{Q},Q),U=U_{Q},Z=Z_{Q}$ and $J=J_{Q}$. Under these definitions, inequality (87) simplifies to:
$\frac{H(K_{A})}{n}\leq\epsilon_{n}+\delta_{n}+I(X;Y|J)+I(U;J|V)-I(U;Z|V).$ (88)
Additionally, the auxiliary random variables $V_{i}=(J^{i-1},Z_{i+1}^{n},F_{1:r})$ and $U_{i}=K_{A}$ imply the Markov chain relationship $(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)$. Assuming that $n$ is sufficiently large, to prove constraint (24) for the random variables $V_{i},J_{i},Z_{i}$ and $F_{1:r}$ as defined above, we proceed as follows:
$I(V;Z)=\frac{1}{n}\sum_{i=1}^{n}I(V_{i};Z_{i})$ $=\frac{1}{n}\sum_{i=1}^{n}I(J^{i-1},F_{1:r};Z_{i}|Z_{i+1}^{n})$ (89)
$=\frac{1}{n}\sum_{i=1}^{n}I(F_{1:r};Z_{i}|Z_{i+1}^{n})+\frac{1}{n}\sum_{i=1}^{n}I(J^{i-1};Z_{i}|Z_{i+1}^{n},F_{1:r})$
=$\overset{(a)}{=}\frac{1}{n}\sum_{i=1}^{n}I(F_{1:r};Z_{i}|Z_{i+1}^{n})+\frac{1}{n}\sum_{i=1}^{n}I(Z_{i+1}^{n};J_{i}|J^{i-1},F_{1:r})$
$\overset{(b)}{\leq}\frac{1}{n}I(Z^{n};F_{1:r})+\frac{1}{n}\sum_{i=1}^{n}I(J^{i-1},Z_{i+1}^{n},F;J_{i})$
$=\frac{1}{n}I(Z^{n};F_{1:r})+I(V;J)=I(V;J)+H(Z)-\frac{1}{n}H(Z^{n}|F_{1:r})$
$\overset{(c)}{\leq}I(V;J)+H(Z)-\Delta+\epsilon^{\prime}_{n},$
where step (a) follows from Marton-Körner identity, (b) is derived using the chain rule, and (c) utilizes the assumption
$\frac{1}{n}H(Z^{n}|F_{1:r})\geq\Delta-\epsilon^{\prime}_{n}$
for a vanishing sequence $\epsilon^{\prime}_{n}$. Thus, for any $J$ and a specific choice of auxiliary random variables $U$ and $V$ that satisfy the Markov chain $(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)$ and inequality (24), by taking the limit as $n\rightarrow\infty$ in equation (88), we obtain:
$S_{\Delta}(X;Y\|Z)\leq I(X;Y|J)+I(U;J|V)-I(U;Z|V).$
Proof of the cardinality bounds are standard and we skip them. The second part of this theorem for the special case of $Z=f(X,Y)$ follows from Lemma 3 (after applying the proper change of variables). ∎
Lemma 3.
Suppose that $Z=g(X)$. For every joint distribution $P_{V,X,Y}$, the maximizer of $I(U;Y|V)-I(U;Z|V)$ over all $P_{U|V,X}$ is achieved when $H(X|U,V,Z)=0$.
Proof of Lemma 3.
Take some arbitrary tuple $P_{U|V,X}$. By the functional representation lemma, we can find a conditional distribution $P_{W|U,V,Z,X}$ such that $I(Z,U,V;W)=0$ and $H(X|Z,U,V,W)=0$. Now, consider the joint distribution $P_{W|U,V,Z,X}P_{Y|X}$. Let $U^{\prime}=(U,W)$. We claim that changing $U$ to $(U,W)$ would not decrease the expression, and moreover $H(X|Z,U^{\prime},V)=0$. This would complete the proof. We need to show that
$I(U;Y|V)-I(U;Z|V)\leq I(U^{\prime};Y|V)-I(U^{\prime};Z|V)$ (90)
Equivalently,
$0\leq I(W;Y|U,V)-I(W;Z|U,V)$ (91)
which holds since $I(W;U,V,Z)=0$. ∎
Appendix C An equivalent form of Conjecture 1
Note that by setting $Z=Z^{\prime}$, $Y=Y^{\prime}$ and replacing $X$ by $(X,X^{\prime})$ in Conjecture 1, we obtain the following equation (as a special case):
$S(XX^{\prime};Y^{\prime}\|Z^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(Y^{\prime}Z^{\prime};X|X^{\prime}),\qquad\forall P_{XX^{\prime}Y^{\prime}Z^{\prime}}.$ (92)
Observe that $S(XX^{\prime};Y^{\prime}\|Z^{\prime})\geq S(XX^{\prime};XY^{\prime}\|XZ^{\prime})$ because the first terminal can publicly reveal $X$ to all parties. Therefore, (92) can be weakened to
$S(XX^{\prime};XY^{\prime}\|XZ^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(Y^{\prime}Z^{\prime};X|X^{\prime}),\qquad\forall P_{XX^{\prime}Y^{\prime}Z^{\prime}}.$ (93)
Therefore, if Conjecture 1 is true, (93) holds. We now claim the reverse.
Theorem 8.
Conjecture 1 holds if and only if (93) holds.
Proof.
We first show that the conjecture holds if (92) holds. The reason is as follows: since $S(X;Y\|Z)=S(Y;X\|Z)$, from (92), we also obtain the symmetric form:
$S(X^{\prime};YY^{\prime}\|Z^{\prime})-S(X^{\prime};Y^{\prime}\|Z^{\prime})\leq I(X^{\prime}Z^{\prime};Y|Y^{\prime}).$ (94)
In particular, (94) implies
$S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(XX^{\prime};Y^{\prime}|Z^{\prime})\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime}).$ (95)
Adding up (92) and (95) we obtain
$S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})\leq I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (96)
Finally, note that
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}|Z^{\prime})$ $=S(X;Y\|Z)-S(XX^{\prime};YY^{\prime}|Z^{\prime})+S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$
$\leq S(X;Y\|Z)-S(X;Y|Z^{\prime})+S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$
$\leq S_{\text{ow}}(XY\rightarrow Z^{\prime}|Z)+S(XX^{\prime};YY^{\prime}|Z^{\prime})-S(X^{\prime};Y^{\prime}|Z^{\prime})$ (97)
$\leq S_{\text{ow}}(XY\rightarrow Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime})$ (98)
where (97) follows from (6), and (98) follows from (96).
It remains to show that (93) implies (92). Equation (92) follows if one can prove the following:
$S(X;Y\|Z)\leq\inf_{\tilde{X}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}S(\tilde{X};Y\|Z)+I(X;YZ|\tilde{X}).$
Let
$\Phi(X;Y\|Z)=\inf_{\tilde{X}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}S(\tilde{X};Y\|Z)+I(X;YZ|\tilde{X}).$ (99)
We use the approach in [14]. To prove $S(X;Y\|Z)\leq\Phi(X;Y\|Z)$, choose an arbitrary $(n,R,\delta)$ code, and let $W_{1}$ and $W_{2}$ be the private randomness used by the two parties, i.e., $H(F_{i}|X^{n},F^{i-1},W_{1})=0$ for odd $i$ and $H(F_{i}|Y^{n},F^{i-1},W_{2})=0$ for even $i$. Then, we have
$\frac{1}{n}H(K_{B})$ $=\frac{1}{n}I(X^{n}F_{1:r}W_{1};K_{B})+\epsilon_{n}$ (100)
$=\frac{1}{n}(I(X^{n}F_{1:r}W_{1};K_{B})-I(K_{B};Z^{n}F_{1:r}))+2\epsilon_{n}$ (101)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r}W_{1};K_{B}\|Z^{n}F_{1:r})+2\epsilon_{n}$ (102)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r}W_{1};Y^{n}F_{1:r}W_{2}\|Z^{n}F_{1:r})+2\epsilon_{n}$ (103)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r-1}W_{1};Y^{n}F_{1:r-1}W_{2}\|Z^{n}F_{1:r-1})+2\epsilon_{n}$ (104)
$\leq\frac{1}{n}\Phi(X^{n}F_{1:r-2}W_{1};Y^{n}F_{1:r-2}W_{2}\|Z^{n}F_{1:r-2})+2\epsilon_{n}$ (105)
[…]
$\leq\frac{1}{n}\Phi(X^{n}W_{1};Y^{n}W_{2}\|Z^{n})+2\epsilon_{n}$ (107)
$=\frac{1}{n}\Phi(X^{n};Y^{n}\|Z^{n})+2\epsilon_{n}$ (108)
$\leq\Phi(X;Y\|Z)+2\epsilon_{n}$ (109)
where $\epsilon_{n}$ converges to zero as $\delta$ converges to zero, (100) follows from the fact that $K_{A}=K_{B}$ with probability $1-\delta$ and $H(K_{A}|X^{n}F_{1:r}W_{1})=0$, (101) follows from the privacy constraint, (102) from property iv in Lemma 4, (103) follows from property iii in Lemma 4 and the fact that $K_{B}$ is a function of $(Y^{n}W_{2}F_{1:r})$, (104)-(107) follow from property ii in Lemma 4 and the fact that $H(F_{i}|X^{n},F^{i-1},W_{1})=0$ for odd $i$ and $H(F_{i}|Y^{n},F^{i-1},W_{2})=0$ for even $i$, (108) follows from property v in Lemma 4, and (109) follows from property i in Lemma 4.
Lemma 4.
Assuming (93), the function $\Phi$ as defined in (99) satisfies the following properties:
  1. i)
    $\Phi(X^{n};Y^{n}\|Z^{n})\leq n\Phi(X;Y\|Z)$ when $P_{X^{n},Y^{n},Z^{n}}=\prod_{i=1}^{n}P_{X_{i},Y_{i},Z_{i}}$ is i.i.d,
  2. ii)
    $\Phi(F,X;F,Y\|F,Z)\leq\Phi(F,X;Y\|Z)$ and $\Phi(F,X;F,Y\|F,Z)\leq\Phi(X;F,Y\|Z)$ for any arbitrary $P_{F,X,Y,Z}$,
  3. iii)
    $\Phi(X;Y^{\prime}\|Z)\leq\Phi(X;Y\|Z)$ when $H(Y^{\prime}|Y)=0$,
  4. iv)
    $\Phi(X;Y\|Z)\geq I(X;Y)-I(Y;Z)$,
  5. v)
    $\Phi(XW_{1};YW_{2}\|Z)=\Phi(X;Y\|Z)$ if $P_{W_{1},W_{2},X,Y,Z}=P_{W_{1}}P_{W_{2}}P_{XYZ}$.
Proof.
Observe that $S(X;Y\|Z)$ similar properties as above properties, i.e., $S(X^{n};Y^{n}\|Z^{n})=nS(X;Y\|Z)$, and $S(F,X;F,Y\|F,Z)\leq S(F,X;Y\|Z)$, etc; see [14].
Proof of property i) Let $\epsilon>0$ be arbitrary. Take some $\tilde{X}$ satisfying $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)$ such that
$S(\tilde{X};Y\|Z)+I(X;Y,Z|\tilde{X})\leq\Phi(X;Y\|Z)+\epsilon.$
Take $n$, i.i.d copies of $(X,Y,Z,\tilde{X})$ and denote it by $(X^{n},Y^{n},Z^{n},\tilde{X}^{n})$. We have:
$\Phi(X^{n};Y^{n}\|Z^{n})\leq S(\tilde{X}^{n};Y^{n}\|Z^{n})+I(X^{n};Y^{n},Z^{n}|\tilde{X}^{n})=n(S(\tilde{X};Y\|Z)+I(X;Y,Z|\tilde{X}))\leq n\Phi(X;Y\|Z)+n\epsilon.$ (110)
Proof of property ii) We first show that $\Phi(F,X;F,Y\|F,Z)\leq\Phi(F,X;Y\|Z)$. Let $\epsilon>0$ be arbitrary. Take some $\tilde{X}$ satisfying $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-(F,X)-\!\!\!\!\minuso\!\!\!\!-(Y,Z)$ such that
$S(\tilde{X};Y\|Z)+I(F,X;Y,Z|\tilde{X})\leq\Phi(F,X;Y\|Z)+\epsilon.$
The Markov chain $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-(F,X)-\!\!\!\!\minuso\!\!\!\!-(Y,Z)$ implies the Markov chain $(F,\tilde{X})-\!\!\!\!\minuso\!\!\!\!-(F,X)-\!\!\!\!\minuso\!\!\!\!-(F,Y,Z)$. Let $X^{\prime}=(F,\tilde{X})$. We have
$\Phi(F,X;F,Y\|F,Z)$ $\leq S(X^{\prime};F,Y\|F,Z)+I(F,X;F,Y,Z|X^{\prime})$
$=S(F,\tilde{X};F,Y\|F,Z)+I(F,X;F,Y,Z|F,\tilde{X})$
$=S(F,\tilde{X};F,Y\|F,Z)+I(F,X;Y,Z|\tilde{X})-I(F;YZ|\tilde{X})$
$\leq S(\tilde{X};Y\|Z)+I(F,X;Y,Z|\tilde{X})$ (111)
$\leq\Phi(F,X;Y\|Z)+\epsilon.$
where (111) follows from the assumption in (93).
Next, we show that $\Phi(F,X;F,Y\|F,Z)\leq\Phi(X;F,Y\|Z)$. Let $\epsilon>0$ be arbitrary. Take some $\tilde{X}$ satisfying $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(F,Y,Z)$ such that
$S(\tilde{X};F,Y\|Z)+I(X;F,Y,Z|\tilde{X})\leq\Phi(X;F,Y\|Z)+\epsilon.$
From $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(F,Y,Z)$, we deduce $(F,\tilde{X})-\!\!\!\!\minuso\!\!\!\!-(F,X)-\!\!\!\!\minuso\!\!\!\!-(F,Y,Z)$ holds. Let $X^{\prime}=(F,\tilde{X})$. Thus, we have:
$\Phi(F,X;F,Y\|F,Z)$ $\leq S(X^{\prime};F,Y\|F,Z)+I(F,X;F,Y,Z|X^{\prime})$
$=S(F,\tilde{X};F,Y\|F,Z)+I(F,X;F,Y,Z|F,\tilde{X})$
$=S(F,\tilde{X};F,Y\|F,Z)+I(X;F,Y,Z|\tilde{X})-I(X;F|\tilde{X})$
$\leq S(\tilde{X};F,Y\|Z)+I(X;F,Y,Z|\tilde{X})$
$\leq\Phi(X;F,Y\|Z)+\epsilon.$
Proof of Property iii) Let $\epsilon>0$ be arbitrary. Take some $\tilde{X}$ satisfying $\tilde{X}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)$ such that
$S(\tilde{X};Y\|Z)+I(X;Y,Z|\tilde{X})\leq\Phi(X;Y\|Z)+\epsilon.$
We have
$\Phi(X;Y^{\prime}\|Z)$ $\leq S(\tilde{X};Y^{\prime}|Z)+I(X;Y^{\prime},Z|\tilde{X})\leq S(\tilde{X};Y|Z)+I(X;Y,Z|\tilde{X})\leq\Phi(X;Y\|Z)+\epsilon.$ (112)
Proof of Property iv) We know that $S(X^{\prime};Y\|Z)\geq I(X^{\prime};Y)-I(Y;Z)$. Thus $S(X^{\prime};Y\|Z)+I(X;Y,Z|X^{\prime})\geq I(X^{\prime};Y)-I(Y;Z)+I(X;Y,Z|X^{\prime})=I(X;Y)-I(Y;Z)+I(X;Z|X^{\prime},Y)\geq I(X;Y)-I(Y;Z).$ This completes the proof.
Proof of Property v) Take some arbitrary $P_{X^{\prime}|X}$ reaching within $\epsilon$ of the infinimum defining $\Phi(X;Y|Z)$. If we define $(W_{1},W_{2})$ independent of $(X^{\prime},X,Y,Z)$, the Markov chain $X^{\prime}-\!\!\!\!\minuso\!\!\!\!-(W_{1},X)-\!\!\!\!\minuso\!\!\!\!-(W_{2},Y,Z)$ will hold. Thus,
$\Phi(X,W_{1};Y,W_{2}\|Z)\leq S(X^{\prime};W2,Y\|Z)+I(W_{1},X;W_{2},Y,Z\|X^{\prime})=S(X^{\prime};Y\|Z)+I(X;Y,Z\|X^{\prime})\leq\Phi(X;Y|Z)+\epsilon.$
This implies $\Phi(X,W_{1};Y,W_{2}\|Z)\leq\Phi(X;Y\|Z)$. Next, to show $\Phi(X;Y\|Z)\leq\Phi(X,W_{1};Y,W_{2}\|Z)$, take some $X^{\prime}$ satisfying $X^{\prime}-\!\!\!\!\minuso\!\!\!\!-(W_{1},X)-\!\!\!\!\minuso\!\!\!\!-(W_{2},Y,Z)$. Then, we have $X^{\prime}-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)$ and
$S(X^{\prime};W_{2},Y\|Z)+I(W_{1},X;W_{2},Y,Z|X^{\prime})\geq S(X^{\prime};Y\|Z)+I(X;Y,Z|X^{\prime})\geq\Phi(X;Y\|Z).$